Systems Staff Meeting Notes 1/11/99

Systems Staff Meeting Notes
January 11, 1998

Present: Rich Bennett, Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Bobbie Meng, Suzy Shaw, Ying Tang
  1. Computer Repair Tech position - An offer is pending, waiting for Personnel to give the word.
  2. Furbys are banned at the National Security Agency.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. There is no distribution list yet for the new Gateways.
  2. Suzanne Kiker's machine is still experiencing problems.
  3. Ashley installed the new machine (Brian's old Gateway) in Debra's office.
  4. Winston's old machine will be swapped out for Bobbie's, Bobbie's will go to Scott & Scott's will replace the Dee One System in the public area of Reference.
  5. We will keep both the 3.2 & the 6.4 IDE drives - they are hard to acquire on the market at this time.
  1. At the last meeting, there was a discussion about a new virus, Picture.exe. At first it had the appearance of a hoax, but it eventually turned up on the McAfee site. The virus sniffs around, looks for an AOL account, checks where you've been & emails the information to China. To remove, you need to do a search for Note.exe, then delete. Bill will send out the information about it to the Liaison list & Debra will archive it to the Liaison section of our web page.
  1. There was a Bell South fiber cut over the weekend outside of Jacksonville. NERDC was slow in getting the notice out. The outage lasted 4-5 hours, but there was no notice until about 30 minutes before it was repaired.
  2. Resource Services wiring - Ashley is wiring it all to a patch panel, then a jumper to the hub. He will train Bobbie & the new Computer Tech in this process. Ashley notes he needs another tool for punching down.
  3. The most recent phone outage occurred during the installation of the Resource Services phones. The box was reset and, as a result, we lost our outside lines.
  4. Special Collections is requesting a "walk-up puddle" (multiple jacks which are activated to allow Internet access) to be located in the Grand Reading Room.
Software Projects Status
  1. Mellon meeting is scheduled for January 26. Winston has been having difficulties with his new NT machine, so he has been unable to work on Mellon lately.
  2. FCLA backups ran last week.
  3. Student Patron Update - initial run complete, second update this week.
  4. Special Collection Imaging Database - John Ingram told Winston that he wants to get rid of the equipment. Bill has okayed this - it will eventually have massive Y2K problems. Get ready to survey out all microchannel type stuff. Their large monitor is black & white - Preservation may have use for it.
  5. Ying is working on updating the Trouble Ticketing System.
  6. PAULA - Bill & Bobbie's machines have been generating error messages. An attempt to remove the HD code located errors between PAULA & the server. Brian says he is working on having PAULA hit the server one time, then have it require a refresh to hit it again.
  7. Newspaper Database - there is not enough information in their database to have it extract the details they want. We can only create a Boolean search, then they can make the fields & fill in.
  8. LASSIE - Brian has conferred with Rich & will set up access for Rich to track his IP's.
  9. The Patron Tracking database for Special Collections is complete.
  10. Bill notes the finding aids that are in Mac format in Special Collections need to be converted. Winston remarks that John Ingram's book has special Greek dialects on his Mac Performa that are unable to be converted or replaced.
  11. SELMA works using certain bitmaps, and the T-sizer is not working as well as usual. Bill suggests getting hold of the T-sizer people for updates.
  12. Brian Burton announces that he & Pam Cenzer are on a search committee for a "weekend John Ashcraft".
  13. CLEO & ILL - all on the same page - need a commonly useable, multi-accessible scratchpad to hold updates & then the SAS system could be removed from use.
  14. Facilities suite is on hold.
  15. ATC is becoming increasingly important.
DP Coordinator
  1. Student schedules are solidified. Will be added to SELMA today.
  2. Bill has not reached Compustat regarding the unreadable yearly tapes.
  1. Ashley Jones will be out Friday, January 15.
  2. Martin Luther King Day - holiday - January 18.
  3. Winston will be out Jan. 25-29.
  4. Debra will be out the afternoon of Jan. 25.
  5. Bill Covey out Jan. 29.
  6. ALA will be Jan. 28 - Feb. 5 in Philadelphia, PA.
  7. Rich will be out to ALA Jan. 29 - Feb. 2.

Last updated February 18, 2000
by Debra Harris